Find A and B so that the graph of Ax+By=-26 contains the points (4,-5) and (1,-11)
Answer by mananth(16946)

Plug the values of the points in the equation Ax+by=-26
we get
4.00 a -5 b = -26 .............1
1 a -11 b = -26 .............2
Eliminate b
multiply (1)by -11
Multiply (2) by 5
-44 a 55 b = 286
5 a + -55 b = -130
Add the two equations
-39 a = 156.000
/ -39
a = -4
plug value of a in (1)
4 a + -5 b = -26
-16 + -5 b = -26
-5 b = -26 + 16
-5 b = -10
b = 2